A conducting square loop is in a uniform magnetic
Answer #1

a) magnitude of induced emf = A*dB/dt

= 0.6^2*(0.05 - 0)/0.2

= 0.09 volts

b) I = induced emf/R

= 0.09/0.045

= 2 A

Direction : clockwise

c) Rate of Thermal energy generated = I^2*R

= 2^2*0.045

= 0.18 A

d) from t = 0 to t = 0.2, emf is constnt and emf = 0.09 volts

from t = 0.2 s to 0.5 s emf = 0


e) induced emf = B*dA/dt

= B*d(L^2)/dt

induced emf = B*2*L*dL/dt

dL/dt = induced emf/(2*B*L)

= 0.09/(2*0.05*0.6)

= 1.5 m/s